OpenStudy (anonymous):

Is this better? Jason has to choose a combination locker. The combination must have 3 numbers. The first and third numbers have to be numbers from 1 to 25. The second number has to be a number from 0 to 4. Find the total number of combinations Jason could choose. Please explain :)

OpenStudy (anonymous):

much better. counting principle. 25 choices for first number, 5 for second and then another 25 for third. multiply \[25\times 5\times 25\] for your answer

OpenStudy (anonymous):

So when it says 1 to 25 , its not the number in between? Cuss thats what some one told me :o & i got it wrong :(

OpenStudy (anonymous):

??

OpenStudy (amistre64):

yes, much better :) we can draw a picture for this but it gets rather larger real quick; the idea tho is tho picture 25 branches to start with, these are all of our options each branch has 5 fingers to it for our second option (0,1,2,3,4) 25*5 = 125 so far each of the 5 tips has 25 other fingers branching out from them; 125 * 25 each = .....

OpenStudy (amistre64):

if my calculator is telling me the truth, id say its 3125

OpenStudy (anonymous):

if i have from 1 to 25 dollars in my pocket my understanding of english is that i could have 1 through 25 dollars, not 2 through 24

OpenStudy (amistre64):

attaching the picture is fine, it helps us to see if you made any typos or not :)